# On learning racing policies with reinforcement learning

### v1

@@ IV-B Track representation
A key aspect of training an effective racing policy is the choice of track representation, which influences learning efficiency and generalization capabilities. Two commonly used approaches for representing track information in RL policies are:
@@
    Track progress representation: The track information is represented in terms of Frenet coordinate $s$, which represents progress along the centerline. This approach ensures that the policy always has precise information about its location on the track. However, unlike the geometric representation, it lacks explicit information about the upcoming track shape, limiting the policy’s ability to generalize across similar sections.
@@ IV-C Action space
Another critical policy design choice is the action space selection, which can directly impact both performance and sim-to-real transfer. We evaluated the effect of control space selection on policy performance and sim-to-real transfer. Using an otherwise identical training configuration, we compared two control architectures: one in which the policy $\pi_{\omega}$ directly controls vehicle wheel speed $\omega_{ref}$ and another in which it modulates the acceleration of the clamped wheel speed $\dot{\omega}_{ref}$. In the simulation, the velocity control policy converged to performance levels comparable to the acceleration control policy, as demonstrated in Figure 4. However, in real-world tests, the wheel speed control-based policy proved overly aggressive, frequently exceeding track boundaries by more than 20 cm and failing to complete 10 laps without crashing. These results underscore the critical importance of choosing a control space that limits abrupt control actions, thereby enhancing the robustness of sim-to-real transfer.
@@ IV-D Actuator modeling
Actuator dynamics is often neglected in vehicle dynamics modeling. In our work, both the powertrain and steering actuators are modeled as first-order systems, as described in Equation (1). To assess the impact of actuator modeling on policy performance, we trained a policy without modeled actuator dynamics, deployed it in the real world, and presented its performance in Table I.
Figure 4: Learning curves of the base policy using $\pi_{\dot{\omega}}$ with $obs$, the policy using $\mathfrak{s}$ for track representation $obs_{\mathfrak{s}}$, and the wheel speed control policy $\pi_{\omega}$.
@@
The policy trained without modeled actuators RL w/o am is significantly slower than the baseline RL policy (see Table I). As illustrated in Figure 5, the policy with modeled actuators initiates steering adjustments earlier, effectively compensating for actuator dynamics, whereas the baseline policy exhibits delayed responses. This finding underscores the importance of incorporating accurate actuator models to improve policy performance and enable more reliable real-world deployment.
@@ IV-E MPC baseline
The MPC controller was implemented with an objective that was as similar to the RL policy as possible, with minimal addition of auxiliary losses to ensure real-time optimization convergence. Furthermore, the dynamics model used in the MPC was identical to the one in the simulation environment, resulting in our fastest implementation of the MPC controller.
@@ IV-F Policy performance
To evaluate our racing policy, we compared its performance against a state-of-the-art MPC approach and an expert human driver—who placed fifth in the Polish Indoor RC Championships.
@@
The RL policy outperformed both the MPC approach and the expert human driver. To the best of our knowledge, this is the first time that the RL policy has surpassed expert human performance in RC car racing. Notably, MPC achieved the lowest $\boldsymbol{E}_{\text{off}}$, while the human driver recorded the highest. This discrepancy may be attributed to the track’s relatively small size compared to the vehicle and the low tire-road friction conditions, which differed significantly from those typically encountered in RC racing competitions. Unlike traditional RC race tracks, which are often carpeted to enhance grip, our test track’s slippery surface may have posed an unexpected challenge for the human driver, limiting their ability to fully exploit the vehicle’s handling capabilities.
The racing lines followed by the human driver and the RL policy — illustrated in the left image of Figure 6 — are similar. Although the RL policy initially takes a tighter arc from $\mathfrak{s}=0m$ to $\mathfrak{s}=5m$, it does not immediately gain time over the human driver but instead positions the vehicle more favorably for the upcoming sections. Between $\mathfrak{s}=5m$ and $\mathfrak{s}=11m$, while both maintain the same longitudinal velocity ($v_{x}$), the RL policy gains approximately 0.14 seconds by adopting a more efficient racing line. In the final segment, from $\mathfrak{s}=11m$ to $\mathfrak{s}=17m$, the policy carries more speed into the corner, securing an additional 0.15-second advantage, resulting in a total lead of nearly 0.3 seconds over the expert human driver for the entire lap.
@@ V Conclusion
Despite these advancements, several challenges remain. While our policy demonstrates superior performance in single-lap racing, future work should focus on adapting to dynamically changing track conditions and incorporating real-time model adaptation. Moreover, further exploration of hybrid learning-based and model-based approaches may yield even greater improvements in performance and safety.

### v2

@@ IV-B Track representation
A key aspect of training an effective racing policy is the choice of track representation, which influences learning efficiency and generalization capabilities. Two commonly used approaches to represent track information in RL policies are:
@@
    Track progress representation: The track information is represented in terms of the Frenet coordinate $s$, which represents progress along the centerline. This approach ensures that the policy always has precise information about its location on the track. However, unlike the geometric representation, it lacks explicit information about the upcoming track shape, limiting the policy’s ability to generalize across similar sections.
@@ IV-C Action space
Another critical policy design choice is the action space selection, which can directly impact both performance and sim-to-real transfer. We evaluated the effect of control space selection on policy performance and sim-to-real transfer. Using an otherwise identical training configuration, we compared two control architectures: one in which the policy $\pi_{\omega}$ directly controls vehicle wheel speed $\omega_{ref}$ and another in which it modulates the acceleration of the clamped wheel speed $\dot{\omega}_{ref}$. In the simulation, the velocity control policy converged to performance levels comparable to the acceleration control policy, as demonstrated in Figure 4. However, in real-world tests, the wheel speed control-based policy proved overly aggressive, frequently exceeding track boundaries by more than 20 cm and failing to complete 10 laps without crashing. These results underscore the critical importance of choosing a control space that limits abrupt control actions, thus enhancing the robustness of sim-to-real transfer.
@@ IV-D Actuator modeling
Actuator dynamics is often neglected in vehicle dynamics modeling. In our work, both the powertrain and the steering actuators are modeled as first-order systems, as described in Equation (1). To assess the impact of actuator modeling on policy performance, we trained a policy without modeled actuator dynamics, deployed it in the real world, and presented its performance in Table I.
Figure 4: Learning curves of the base policy using $\pi_{\dot{\omega}}$ with obsobs, the policy using $\mathfrak{s}$ for track representation $obs_{\mathfrak{s}}$, and the wheel speed control policy $\pi_{\omega}$.
@@
The policy trained without modeled actuators RL w/o am is significantly slower than the baseline RL policy (see Table I). As illustrated in Figure 5, the policy with modeled actuators initiates steering adjustments earlier, effectively compensating for actuator dynamics, while the baseline policy exhibits delayed responses. This finding underscores the importance of incorporating accurate actuator models to improve policy performance and enable more reliable real-world deployment.
@@ IV-E MPC baseline
The MPC controller was implemented with an objective that was as similar to the RL policy as possible, with minimal addition of auxiliary losses to ensure real-time optimization convergence. Furthermore, the dynamics model used in the MPC was identical to the one in the simulation environment, resulting in our best-performing MPC implementation.
@@ IV-F Policy performance
To evaluate our racing policy, we compared its performance against a state-of-the-art MPC approach and an expert human driver, who placed fifth in the Polish Indoor RC Championships.
@@
The RL policy outperformed the MPC baseline in both fastest-lap and average-lap times. We attribute RL ’s advantage, at least in part, to MPC ’s finite prediction horizon, which biases control toward locally optimal actions rather than lap-time–optimal trajectories. Moreover, the RL policy surpassed the expert human driver across all evaluated metrics.To the best of our knowledge, this is the first time that the RL policy has exceeded expert human performance in RC car racing. Notably, MPC achieved the lowest $\boldsymbol{E}_{\text{off}}$, while the human driver recorded the highest. This discrepancy may be attributed to the track’s relatively small size compared to the vehicle and the low tire-road friction conditions, which differed significantly from those typically encountered in RC racing competitions. Unlike traditional RC race tracks, which are often carpeted to enhance grip, our test track’s slippery surface may have posed an unexpected challenge for the human driver, limiting their ability to fully exploit the vehicle’s handling capabilities.
The racing lines followed by the human driver and the RL policy — illustrated in the left image of Figure 6 — are similar. Although the RL policy initially takes a tighter arc from $\mathfrak{s}=0~m$ to $\mathfrak{s}=5~m$, it does not immediately gain time over the human driver, but instead positions the vehicle more favorably for the upcoming sections. Between $\mathfrak{s}=5~m$ and $\mathfrak{s}=11~m$, while both maintain the same longitudinal velocity ($v_{x}$), the RL policy gains approximately 0.14 seconds by adopting a more efficient racing line. In the final segment, from $\mathfrak{s}=11~m$ to $\mathfrak{s}=17~m$, the policy carries more speed into the corner, securing an additional 0.15-second advantage, resulting in a total lead of nearly 0.3 seconds over the expert human driver for the entire lap.
@@ V Conclusion
Despite these advancements, several challenges remain. Although our policy demonstrates superior performance in single-lap racing, future work should focus on adapting to dynamically changing track conditions and incorporating real-time model adaptation. Moreover, further exploration of hybrid learning-based and model-based approaches may yield even greater improvements in performance and safety.